OpenStudy (anonymous):

Simplify the expression. √-4/(5+2i)-(3+4i) The i is an imaginary #.

OpenStudy (anonymous):

\[\sqrt{-4}/(3-2i)\]

OpenStudy (anonymous):

sqrt of -4 is generally an impossible or undefined component of math....also an imaginary number

OpenStudy (anonymous):

That's wrong. The choices are a) 3+i/10, b) 3-i/10, c) 3+i/8, d)3-i/8

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So I'm half right, and half in the obliquesphere...or you put the wrong eq up there...

jhonyy9 (jhonyy9):

- all terms are inside sqrt ?

OpenStudy (anonymous):

Thats what I'm wondering... is it the 4 thats neg, or the sqrt of 4?

OpenStudy (anonymous):

and is it just the four under sqrt, or all of it?

jhonyy9 (jhonyy9):

denominator is 5+2i or this all without -4 ?

OpenStudy (anonymous):

it's √-4. So I know that it's impossible to find the square root of a negative # that's why it turns into an imaginary number. It would be i√4. Then you find the prime factors of 4 which are 2&2. Group the pair. So it turns into 2i. I just don't know how to do the rest and get the correct answer.

OpenStudy (anonymous):

It's √-4 over (5+2i)-(3+4i)

OpenStudy (anonymous):

3+(i/10)

jhonyy9 (jhonyy9):

so but inside sqrt is just 4 or the denominator too ?

jhonyy9 (jhonyy9):

2i i ------ = ------ 2-2i 1-i

OpenStudy (anonymous):

√-4/(5+2i)-(3+4i) (2i)/(3-2i) 3+i/8 (oh sorry lol)

jhonyy9 (jhonyy9):

how do you like this ?

OpenStudy (anonymous):

/facepalm I hated that question.

OpenStudy (anonymous):

"Please sir, may I have another?"

OpenStudy (anonymous):

(in another thread)

jhonyy9 (jhonyy9):

denominator is (5+2i) -(3+4i) ?

jhonyy9 (jhonyy9):

Dyiliq check it please 5-3 is not equal 3

OpenStudy (anonymous):

√-4/(5+2i)-(3+4i) (2i)/(2-2i) 3+i/8 woops

jhonyy9 (jhonyy9):

why ? 2i i ------- = ------- 2(1+i) 1+i

jhonyy9 (jhonyy9):

sorry i/(1-i)

jhonyy9 (jhonyy9):

i(1+i)/(1-(-1)) = (i - 1)/2

OpenStudy (anonymous):

invert 3+i/8 = i/2 (2i)/(3-2i) = i/2 √-4/(5+2i)-(3+4i) = i/2

OpenStudy (anonymous):

sorry again lolz (2i)/('2'-2i)

OpenStudy (anonymous):

and 3+i/8= (-i)/(2)

OpenStudy (anonymous):

(2i)/(2-2i) = (-i)/2 √-4/(5+2i)-(3+4i) = (-i)/2